### v1

@@ 6.1.1 Stratified Results for Regression
In the previous sections, we presented the aggregated results across all 137 regression datasets. We now extend the analysis by stratifying the results according to the dataset metrics introduced in Section 5.1, allowing us to capture more nuanced behaviors of the regressors.
Figure 7 presents the stratified results, where each plot corresponds to one metric and divides the datasets into three bins representing the terciles. When multiple datasets share a metric value that lies at the boundary between two bins, those datasets are included in both bins.
When considering the number of samples, we observe an improvement in the performance of k-NN, as expected, and of Decision Trees as the number of samples increases. Conversely, Symbolic Regression exhibits a deterioration in relative performance. In this regime, the selected SR library (PySR) enables batching to make the symbolic model search more efficient, as suggested in the official documentation [^2] https://astroautomata.com/PySR/v1.5.9/api. Despite this, Symbolic Regression appears to struggle in such settings, possibly due to inefficiencies in the symbolic search process when handling large and potentially noisy datasets. This behavior is also reflected in the T2 and T3 metrics, where the relative performance of SR decreases at lower values of these metrics.
The opposite trend is observed when considering the number of features. Surprisingly, SR seems relatively insensitive to higher dimensionality, contrary to common expectations. This may be explained by the feature selection step performed by PySR, which relies on a random forest to reduce dimensionality before model search.
As expected, in high-dimensional spaces, instance-based methods such as k-NN suffer from the curse of dimensionality. PR+LASSO is also affected, though to a lesser extent, as the polynomial expansion increases the number of features, thereby magnifying the impact of high dimensionality.
When analyzing the linearity of the problem using L1, we find that when datasets are well described by a linear model (i.e., low L1 values), nearly all regressors, except for DTs and k-NN, achieve comparable performance. As non-linearity increases, methods capable of modeling non-linear relationships, such as SR, EBM, and IGANN, outperform the others. PR+LASSO, despite being a linear combination of polynomial terms, appears less sensitive to increasing non-linearity, likely due to its flexible basis expansion.
For the local noise metric, S3, derived from the error of a 1-NN regressor, we do not observe strong or unexpected patterns, except in the highest noise regime, where k-NN surprisingly does not perform the worst. This may indicate that other regressors are similarly affected by noise, reducing relative differences.
In datasets containing a feature that is highly correlated with the target (higher values for metric $\text{C1}_{\text{r}}$), all regressors perform more similarly. When the features are weakly correlated, we observe the expected distribution of model performances, with k-NN and Decision Trees outperforming linear models.
Finally, we note that the various linear regression variants perform comparably overall, except for LASSO, which shows advantages in high-dimensional settings (metrics T2, T3, and number of columns), confirming its intended regularization behavior.
6.2 Out-Of-Sample Regression
The previous sections presented the results obtained in the in-sample setting. To further deepen our analysis, we now examine the results derived from the out-of-sample simulations, as described in Section 4.2. The results reveal some notable behavioral differences compared to the in-sample setting. On the right-hand side of Figure 5, we report the aggregated rankings in the out-of-sample regime. Here, we can observe that Linear Regression and standard Decision Trees are adversely impacted by the out-of-sample setting, to the advantage of GLM, LASSO, and k-NN.
To better understand the origin of these differences, we examine the stratified results reported in Figure 8. The overall pattern remains consistent across most stratifications, with some deviations that complement the trends discussed in Section 6.1.1. For instance, k-NN tends to perform relatively better in the low-sample regime, while SR appears to be more adversely affected. When the number of features is small, Polynomial Regression, LASSO, and GLM show relative improvements. Conversely, k-NN is less negatively impacted by higher dimensionality, in contrast to its behavior under in-sample conditions. These tendencies are also reflected in metrics T2 and T3. Furthermore, GLM and LASSO tend to outperform Linear Regression as the linearity of the dataset decreases (metric L1). k-NN appears less affected by local noise (metric S3), while both GLM and LASSO achieve better performance when a feature exhibits strong correlation with the target variable (metric $\text{C1}_{\text{r}}$). In the same regime, Symbolic Regression faces a relative decay of its predictive performance.
Figure 8: Stratified results for the regression task (out-of-sample).
6.2.1 Training Time for Regression
Figure 9: Training times, measured in seconds, for the regression models. The horizontal dashed line represents the maximum allowed time for each method (300 seconds).
Figure 9 presents the training time of the evaluated models, a crucial aspect in scenarios requiring responsive model updates or rapid retraining.
Among the analyzed methods, IGANN exhibits the longest execution times, followed by EBM. Since IGANN relies on neural networks to compute shape functions, it would likely benefit from GPU acceleration. PR+LASSO shows the next highest training time, largely due to the expanded feature space created by the polynomial transformation. Interestingly, even GLM demonstrates relatively long training times despite its comparatively weak predictive performance.
Subsequently, k-NN, DT, LASSO, and Linear Regression show shorter training times. We emphasize the high variability in training times across methods, which highlights substantial differences in computational efficiency among interpretable models. A general trend also emerges: more expressive methods, those capable of representing richer classes of functions, tend to require significantly longer training times. Aside from GLM, methods such as EBM, IGANN, and PR+LASSO offer better predictive performance but demand one to two orders of magnitude more computation.
In Figure 9, Symbolic Regression is omitted since its execution time consistently reached the maximum threshold. Symbolic Regression requires extensive exploration of the functional space, a process that becomes particularly demanding for datasets with many samples, as each candidate expression must eventually be evaluated on the entire dataset (even with batching). This behavior also explains the underperformance of Symbolic Regression observed in the stratified analysis by sample size presented in Section 6.1.1.
7 Classification Results
7.1 Classification Performance (In-Sample)
Figure 10: Ranking comparison for the classification task between in-sample and out-of-sample settings.
For the classification task, we adopt the same analysis procedure used for regression, this time employing the F1 metric. Figure 10 reports the aggregated rankings computed across all 79 classification datasets; with the left-hand side displaying the in-sample results.
Across all classification datasets, three distinct groups of classifiers can be identified, although the separation between them is less pronounced than in the regression setting. EBM achieves the best performance by a clear margin, a trend that will also be observed in subsequent analyses. The second group comprises all remaining classifiers except Naive Bayes, which consistently performs the worst by a substantial margin.
It is noteworthy that traditional Decision Trees and GOSDT exhibit comparable predictive performance, with a slight advantage in favor of the former. Given that both methods produce similar model structures (i.e., decision trees), their interpretability can be directly compared by analyzing the size of the resulting trees. Figure 11 shows the distribution of the number of nodes in the retrieved models across all datasets. As expected, GOSDT yields significantly sparser trees, indicating more compact and interpretable models.
Figure 11: Comparison of model complexity, measured by tree size, between traditional Decision Trees and GOSDT.
7.1.1 Stratified Results for Classification
@@
As in Section 6.1.1, we stratify the classification results according to the dataset metrics introduced in Section 5.1. Figure 12 presents the stratified results for each metric.
Overall, the considered dataset metrics provide limited stratification of classifier performance (in contrast to the regression case), largely due to the smaller number of classification datasets (79 vs. 137). When stratifying into terciles or quartiles, the resulting strata contained too few datasets to yield meaningful comparisons. For this reason, we opted to stratify the classification datasets into halves, which provides a sample size per stratum comparable to that used in the regression analysis.
When considering the structural properties of the datasets (i.e., the number of rows, columns, and their ratio T2), we observe that the methods tend to perform similarly on datasets with a small number of samples or features. For datasets with a larger number of samples, Naive Bayes and GOSDT tend to perform worse, while k-NN and IGANN perform better, as expected. As the number of features increases, IGANN shows improved performance, which is expected given its neural network foundation, while decision-tree-based models, especially GOSDT, exhibit a decline in accuracy.
When examining the ratios T2 and T3, we observe the struggles of k-NN, already noticed in high dimensional feature space. This is consistent with its instance-based nature: as dimensionality grows, the feature space becomes increasingly sparse, and substantially more samples are needed to maintain good generalization. It is also noteworthy that Logistic Regression remains relatively robust both in low-sample settings and in high-dimensional scenarios.
For the dataset separability metric (F1), Logistic Regression and IGANN perform particularly well on highly separable datasets, with their performance declining as separability decreases. For the N3 metric (noisiness assessed via 1-NN error), classifiers tend to behave similarly when noise is high, suggesting noise reduces relative performance differences. Class entropy ($\text{C1}_{\text{c}}$) reveals that GOSDT is particularly sensitive to class imbalance. For the local class-entropy metric (E5), no strong overall trend is visible beyond the weaker performance of decision trees at lower E5 values.
7.2 Out-Of-Sample Classification
On the right-hand side of Figure 10, we report the aggregated results for the out-of-sample regime. It is worth noting that the out-of-sample setting tends to increase performance variability thus reducing the relative differences across methods.
The stratified results, shown in Figure 13, display trends consistent with those observed in the in-sample analysis. Two remarkable deviations emerge from this comparison. First, when considering the number of features, Logistic Regression exhibits a relative improvement as the dimensionality increases. Second, the sensitivity of k-NN to structural metrics is significantly exacerbated in the out-of-sample setting. Specifically, k-NN shows markedly lower performance on low-sample datasets, high-dimensional datasets, and datasets characterized by an unfavorable ratio of rows to columns. This pattern strongly suggests that k-NN is particularly vulnerable when there are insufficient out-of-sample examples to adequately describe or tile the feature space, leading to less reliable neighborhood identification. Furthermore, when assessing the feature separability (metric F1), k-NN tends to underperform for lower F1 values, that is, when a specific feature is highly effective at discriminating between classes, indicating that its reliance on local proximity cannot represent simple, global linear separators that other models, like Logistic Regression, exploit effectively.
Figure 13: Stratified results for the classification task (out-of-sample).
7.2.1 Training Time for Classification
@@
Figure 14: Training times, measured in seconds, for the classification models. The horizontal dashed line represents the maximum allowed time for each method (300 seconds).
The training time evaluation for the classification task follows the same procedure as described for regression. Figure 14 presents the training times of the classification models.
As observed in the regression task, most of the observed results align with expectations. IGANN exhibits the longest training time, often approaching the maximum allowed duration of 300 seconds during final training. GOSDT shows the second-highest training time, approximately two orders of magnitude greater than that of ordinary Decision Trees, which are among the fastest methods. EBM and, unexpectedly, Logistic Regression follow, while k-NN, Decision Trees, and Naive Bayes achieve the shortest training times overall.
Notably, Logistic Regression, despite sharing the same linear formulation, trains slower than Linear Regression in the regression task. This difference arises because Logistic Regression relies on gradient descent optimization, whereas Linear Regression uses the closed-form Ordinary Least Squares solution.
Moreover, an interesting trade-off emerges between interpretability and training time: more interpretable or sparse models, such as IGANN and GOSDT, tend to require longer training times compared to methods like EBM and Decision Trees. This additional computational cost arises from the need to search for sparser solutions, as in GOSDT, or to optimize smoother shape functions, as in IGANN.
@@ 8 Discussion
Our results can be contextualized by examining previous findings in the literature. The work of kraus2023interpretable, who introduced IGANN, also utilized a stratified analysis on PMLB datasets (categorized by sample size and dimensionality). They showed that IGANN could outperform other interpretable models, including EBM, especially on small or low-dimensional data. This contrasts with our findings that EBM often exhibits stronger predictive performance, a difference that we can attribute to the limited hyperparameter search space used in that study for EBM (where the number of possible interactions was set to zero). While this can favor interpretability, it also potentially affects predictive performance. We also observe that, coherently with that study, the execution time of IGANN is significantly worse compared to EBM and other competing models. Our classification results align with theirs, showing EBM generally outperforms IGANN in that task.
kruschel2025challenging evaluated a set of interpretable models on 20 datasets, noting that while black-box models performed the best in 14 out of 20 cases, the performance gap with strong interpretable models (especially from the GAM family) was small. They also attempted to measure model interpretability by using some criteria introduced by sudjianto2021designing, which yielded counter-intuitive results. For instance, they place Decision Trees on a low position on their interpretability scale, despite their widespread acceptance as being highly interpretable. This ambiguity around objective interpretability metrics supports our decision to avoid comparing interpretability, with the only exceptions being decision-tree-based and linear models, since they share the same model structure.
Consistent with the findings of la2021contemporary and wilstrup2021symbolic, our evaluation confirms that Symbolic Regression consistently outperforms traditional Linear Regression models. wilstrup2021symbolic, who tested SR on 48 PMLB datasets, specifically concluded that SR models generalize better to out-of-sample data than Linear Regression, LASSO, and Decision Trees. This evidence aligns with our observation that SR captures non-linear relationships that simple linear models miss, making it a stronger alternative to LR.
Comparisons involving traditional classification families also appear in wainer2016comparison, who analyzed 14 algorithm families on 115 datasets. They found that k-NN generally outperforms Naive Bayes, with GLMNET (related to our Logistic Regression/LASSO models) positioning between the two. This provides a reference point for the ranking of the simpler models in our study, which is consistent with their findings.
Finally, the work of scholz_comparison_2021 provides a relevant contrast due to their stratified analysis on synthetic data based on the structural properties of the data (number of rows and columns) and relations with the target variable. They found that no classifier always dominates and that performance varies across scenarios. Unlike our findings, they concluded that Logistic Regression is fast but not competitive in predictive accuracy; we observed that Logistic Regression remains a competitive option in specific scenarios, such as those involving high dimensionality, scarce data, or the presence of a highly discriminative feature. Consistent with our results, they found that LASSO-type regularization often does not improve performance enough to be a top choice and that individual Decision Tree models are generally not among the strongest performers. A notable divergence, however, concerns k-NN. scholz_comparison_2021 found that k-NN surprisingly achieved a high rank in small, high-dimensional synthetic datasets. In contrast, our analysis on real-world data highlights this specific regime as a weakness for k-NN, particularly in out-of-sample settings where the feature space is not sufficiently tiled by training examples.

### v2

@@ 5.2 Meta-Features Effects on Regressors
To understand the drivers of model performance, we analyze the Spearman correlation between the dataset meta-features and the models’ predictive scores (Figure 8(a)). Only the significant values are colored.
Several strong correlations emerge:
  • Target Correlation ($\text{C1}_{\text{r}}$): Strong positive correlation with performance, especially for linear-based models.
  • Non-linearity (L1): Negatively impacts linear models, while SR, GAMs, and decision trees demonstrate robustness and model non-linearity effectively.
  • Smoothness/Sampling Quality (S3): Negatively affects all methods.
  • Dimensionality and Sparsity: Negatively impacts EBM, DT, IGANN, PR+LASSO, and severely impacts k-NN. Linear models and SR remain relatively robust.
To capture more nuanced behaviors, we stratified the regression datasets into terciles based on their meta-features. Figure 9 presents a selection of these results, with the remainder detailed in Appendix E. The rankings within these strata reveal behaviors largely consistent with theoretical expectations. For instance, k-NN is heavily penalized by data sparsity, whereas LASSO benefits from it. Furthermore, as expected, linear models underperform in non-linear scenarios (L1), where SR conversely excels. At the same time, non-parametric models, such as decision trees and k-NN, underperform in highly linear settings.
Interestingly, while SR performs exceptionally well in low-sample settings, we observe a surprising relative degradation in performance as the sample size increases. We rule out training time constraints as the primary cause for this trend, as confirmed by an ablation study detailed in Appendix F. Instead, we note that the resulting equations remain remarkably small even with larger sample sizes. This suggests that the relative underperformance is likely driven by the strict parsimony of the model selection criterion, which may overly penalize complexity and prevent SR from capturing finer patterns in large datasets. Conversely, SR demonstrates unexpected robustness in high-dimensional settings. This resilience to a large number of features is non-trivial; it demonstrates that the initial feature selection step utilized by PySR effectively mitigates the combinatorial explosion that has traditionally been considered one of the most severe limitations of this approach.
Furthermore, when a feature is highly correlated with the target variable, all methods tend to perform similarly. Finally, target noise (S3) does not appear to differentiate model performance, aside from an expected detrimental effect on k-NN, suggesting that noise degrades the performance of all methods relatively uniformly.
5.3 Investigating the Classification Performance Landscape
Figure 11: PCA biplot of the dataset meta-features projected onto the first two principal components. Individual datasets are color-coded by the winning algorithm.
Figure 12: Hierarchical clustering of method performance profiles across classification datasets.
In Figure 8(b), we observe that classifier performance moderately correlates with certain meta-features, particularly those related to dimensionality and noise, broadly consistent with patterns seen in regression. An additional effect is observed for GOSDT, which exhibits a distinct sensitivity to class imbalance.
Unlike regression, however, stratifying the classification datasets fails to produce any consistent or interpretable performance hierarchy. Figure 10 illustrates an example of this meta-feature-induced stratification. To nearly match the sample sizes used in the regression analysis, we divide the datasets into halves rather than terciles. The only distinct pattern emerges when a highly discriminative feature (the F1 meta-feature) is present: under these conditions, Logistic Regression improves its relative standing, whereas traditional Decision Trees experience a relative decline in performance. Stratifications based on the remaining meta-features (detailed in Appendix E) do not reveal any significant patterns.
Given the absence of a clear hierarchy, even after stratification, we investigate whether underlying structures or clusters exist that our previous aggregated analyses might have missed.
First, we project the meta-features into a reduced space using Principal Component Analysis (PCA). The resulting biplot (Figure 11), which captures 58% of the variance, plots the datasets on the first two principal components, colored by the winning algorithm. This visualization reveals no clear clustering of winning models in the meta-feature space, indicating that even linear combinations of these meta-features fail to meaningfully separate the winning models, indicating that the meta-feature space does not encode a usable structure for model selection.
Next, to determine if the performance profiles themselves naturally cluster, we applied hierarchical clustering (using Euclidean distance and Ward’s method) directly to the methods’ performances across all classification datasets. This yielded a poor overall silhouette score of 0.22 (Figure 12). When we forced the algorithm to find a number of clusters equal to the number of methods, the silhouette score decreased further to 0.05. For reference, the silhouette score computed using this same methodology for the regression tasks was 0.54 and 0.30, respectively. These metrics confirm that distinct performance clusters do not exist in this classification benchmark; rather, model successes and failures are highly interleaved across datasets and do not align with the selected meta-features.
To ensure this behavior was not merely an artifact of the discrete nature of the F1 score, we confirm that this lack of hierarchy persists even when evaluating the models using continuous proper scoring rules, such as the Brier score. We include these results in Appendix G.
Overall, we do not find any evidence of a stable or recoverable performance hierarchy in classification. Taken together, these results indicate that classifier performance cannot be organized into a low-dimensional or meta-feature-driven structure, and that current meta-features are insufficient for reliable characterization.
5.4 Operationalizing Algorithm Selection and Dataset Difficulty
((a)) Regression (in-sample)
((b)) Regression (out-of-sample)
Figure 13: $R^{2}$ scores of the meta-models predicting the $\Delta$ performance between pairs of models in regression (a) in-sample and (b) out-of-sample.
((a)) Regression (in-sample)
((b)) Regression (out-of-sample)
Figure 14: Random Forest feature importances for predicting dataset difficulty in regression (a) in-sample and (b) out-of-sample.
Given the clear contrast between the structured behavior of regression models and the highly interleaved nature of classification performance, a natural next step is to test whether we can leverage these insights predictively. To this aim, we evaluate whether meta-features can be operationalized for algorithm selection and dataset difficulty estimation.
To answer whether we can operationalize algorithm selection using meta-features, we trained a Random Forest meta-model (evaluated via 5-fold cross-validation) to predict the pairwise difference in performance between methods $i$ and $j$, defined as:
$\Delta_{i,j}=S_{i}-S_{j}$ (14)
where $S$ represents the $R^{2}$ or F1 score on a dataset.
In regression (Figure 13(a)), the meta-model achieves consistently positive $R^{2}$ values, effectively capturing differences in inductive biases. The only exceptions are pairs of highly similar methods, such as EBM versus IGANN or closely related linear models. When evaluating out-of-sample behavior (Figure 13(b)), predictive performance decreases, indicating that meta-features are less reliable for anticipating generalization. While the overall structure remains similar to the in-sample setting, the discrepancies are amplified. In particular, the relative performance between certain model families, such as linear and additive methods, becomes difficult to predict.
We further explored what makes a dataset inherently “hard” by training a meta-model to predict the average performance across methods on a given dataset. For regression in-sample (Figure 14(a)), the meta-model achieves strong performance ($R^{2}=0.64$), with S3, $\text{C1}_{\text{r}}$, and L1 as the dominant features, while dimensionality-related features play a minor role. If we look at the feature importance for out-of-sample regression (Figure 14(b)), we still notice that meta-features related to dimensionality are less important. The only difference is that the linearity of the problem seems to have a larger impact on the difficulty of the dataset.
In contrast, and as expected, the same analyses on classification tasks show that the meta-model rarely achieves positive $R^{2}$ values, indicating that meta-features do not provide reliable guidance for algorithm selection. Similarly, dataset difficulty is only weakly captured ($R^{2}=0.28$), suggesting that classification performance cannot be effectively characterized using the available meta-features. Detailed results and figures are provided in Appendix G.
5.5 Structural Interpretability and Model Complexity
((a)) Linear Models
((b)) Decision Tree Models
Figure 15: Comparison of model complexity measured by (a) the number of non-zero linear terms, between Linear Regression, LASSO, and Polynomial Regression with LASSO; and by (b) tree size, between traditional Decision Trees and GOSDT.
Having established the predictive performance landscape and the role of dataset meta-features, we now shift our focus to the operational properties of the models. In IML, predictive accuracy is only one facet of utility; the structural complexity of the resulting model is equally critical.
Quantifying interpretability is inherently architecture-specific, precluding a direct, one-to-one comparison across heterogeneous models (e.g., k-NN versus Symbolic Regression). However, by examining models within the same structural families, we can objectively assess their relative complexities.
Within the family of linear models, interpretability can be mathematically approximated by analyzing the sparsity of the resulting equations. To quantify this, we apply a numerical threshold, treating any coefficient with an absolute value below $10^{-6}$ as zero. Figure 15(a) illustrates the distribution of the remaining non-zero coefficients. The results demonstrate that LASSO effectively minimizes model complexity compared to standard Linear Regression. Conversely, PR+LASSO exhibits a marked increase in the number of terms due to the polynomial expansion of the feature space. Beyond the sheer number of terms, the inclusion of higher-order transformations and interaction effects makes Polynomial Regression inherently less transparent than purely linear models. Nevertheless, this added structural complexity allows the model to capture non-linear relationships that strict linear formulations cannot fundamentally represent. PR+LASSO thus exemplifies a clear performance-interpretability trade-off: increased expressive power can yield improved predictive accuracy but generates elaborate symbolic representations that require significantly greater effort to interpret.
A similar structural comparison can be made within tree-based models. Traditional Decision Trees and GOSDT exhibit comparable predictive performance, with a slight advantage favoring the former. Because both methods yield identical model structures (decision trees), their interpretability can be directly compared via the number of resulting nodes. Figure 15(b) plots the distribution of node counts across all datasets. As anticipated, GOSDT produces significantly sparser trees, delivering more compact and therefore more easily interpretable models.
5.6 Computational Efficiency and Training Time
((a)) Regression Training Time
((b)) Classification Training Time
@@ 5.6 Computational Efficiency and Training Time
Figure 16: Training times, measured in seconds, for (a) the regression models and (b) the classification models. The horizontal dashed line represents the maximum allowed time for each method (300 seconds).
Beyond structural complexity, training time remains a critical operational constraint, particularly in environments requiring rapid retraining or responsive model updates.
Figure 16(a) details the training times for the regression models. A high variance in execution time is evident across the methods, underscoring substantial differences in computational efficiency among interpretable models. IGANN requires the longest execution times, likely due to its reliance on neural networks to compute shape functions, a process that would heavily benefit from GPU acceleration. PR+LASSO is the next slowest, primarily burdened by the expanded feature space created during polynomial transformation. Interestingly, GLM demonstrates relatively long training times despite its comparatively weak predictive performance. At the other end of the spectrum, k-NN, Decision Trees, LASSO, and Linear Regression train rapidly.
Symbolic Regression was omitted from Figure 16(a) because its execution time was intentionally set to reach the maximum allowable threshold. This extended duration was necessary to accommodate the extensive exploration of the functional space that SR requires. Because every candidate expression must eventually be evaluated across all samples, this exploration becomes exceptionally computationally demanding on larger datasets. We confirmed, however, that this time constraint did not artificially degrade performance on larger datasets; supplementary tests on a random subset showed that extending the training duration did not improve predictive accuracy. Additional details are provided in Appendix F.
The classification training times, presented in Figure 16(b), largely mirror the expectations set by the regression task. IGANN again exhibits the longest training durations. GOSDT ranks second, taking approximately two orders of magnitude longer than standard Decision Trees, which remain among the fastest methods overall. EBM and Logistic Regression follow in duration, while k-NN, Decision Trees, and Naive Bayes are highly efficient. It is worth noting that Logistic Regression trains slower than its regression counterpart (Linear Regression) because it relies on iterative gradient descent optimization, whereas Linear Regression utilizes a closed-form Ordinary Least Squares solution.
Ultimately, these timing metrics reveal a pervasive trade-off spanning both regression and classification tasks. More expressive methods capable of representing richer function classes (e.g., EBM, IGANN, PR+LASSO) require one to two orders of magnitude more computation. Furthermore, algorithms explicitly designed to search for sparser, more interpretable solutions (such as GOSDT) or smoother shape functions (like IGANN) incur a heavy computational penalty compared to their standard, less constrained counterparts.
@@ 6 Discussion
To contextualize our results, we first examine them alongside existing comparative benchmarks and the broader literature on data complexity. Following this, we explore the implications of our meta-learning analysis for algorithm selection. We conclude by providing practical guidelines for practitioners seeking to implement interpretable methods.
6.1 Contextualization within Comparative Benchmarks
@@
When examining specific model performances, our results both align with and challenge existing literature. The work of kraus2023interpretable, who introduced IGANN, used a stratified analysis on PMLB datasets and showed that IGANN could outperform other interpretable models, including EBM, especially on small or low-dimensional data. This contrasts with our findings that EBM often exhibits stronger predictive performance, a difference we attribute to the limited hyperparameter search space used in their study for EBM (where the number of possible interactions was set to zero). However, our classification results and observations regarding IGANN’s significantly higher execution times are entirely consistent with their findings.
Consistent with the findings of la2021contemporary and wilstrup2021symbolic, our evaluation confirms that Symbolic Regression consistently outperforms traditional Linear Regression models. wilstrup2021symbolic concluded that SR models generalize better to out-of-sample data than Linear Regression, LASSO, and Decision Trees. This aligns with our observation that SR captures non-linear relationships that simple linear models miss. Expanding upon these prior studies, our results highlight two critical behaviors of SR regarding data dimensionality and sample size. First, SR demonstrates remarkable robustness in high-dimensional settings. This is a non-trivial outcome, as it emphasizes that modern implementations utilizing initial feature selection steps, such as PySR, successfully mitigate the combinatorial explosion that has historically been the primary limitation of symbolic search. Second, contrary to its strength with many features, we observe a surprising degradation in SR’s relative performance as the number of samples increases. As discussed in our results, this underperformance is not driven by training time bottlenecks over large datasets, but rather stems from the strict parsimony of the model selection criterion, which overly penalizes complexity and limits the equation’s ability to capture subtle patterns in massive datasets.
Comparisons involving traditional classification families also align with wainer2016comparison, who found that k-NN generally outperforms Naive Bayes, with penalized linear models positioning between the two. Finally, scholz_comparison_2021 stratified synthetic data and found that no classifier consistently dominates in all settings. While we agree that LASSO-type regularization does not always guarantee top-tier performance and that individual Decision Trees are rarely the strongest performers, we diverge regarding k-NN. Although scholz_comparison_2021 found that k-NN achieved high ranks in small, high-dimensional synthetic datasets, our analysis on real-world data highlights this specific regime as a weakness for k-NN, particularly in out-of-sample settings where the feature space is not sufficiently tiled by training examples.
6.2 Algorithm Selection and the Predictability Gap
A primary contribution of this study is the meta-analysis of algorithm performance, which directly addresses the classic Algorithm Selection Problem (ASP) [khan2020literature]. The ASP posits that no single algorithm is universally superior, a concept formalized by the "No Free Lunch" theorem [wolpert2002no, vanschoren2018meta], and seeks to map measurable characteristics of datasets to expected algorithm performance. Interestingly, while the vast majority of meta-learning literature has historically focused on classification tasks [khan2020literature], our meta-models reveal a profound “predictability gap” between regression and classification.
For regression, algorithm performance and pairwise superiority are highly predictable from dataset meta-features. In contrast, classification performance proves highly idiosyncratic, with meta-features failing to reliably predict task hardness or pairwise differences between methods. Across all analyses, we find no evidence of a stable or recoverable performance hierarchy in classification. Neither stratification, dimensionality reduction, nor clustering reveals any consistent structure, indicating that current meta-features are insufficient for reliable characterization or prediction.
This suggests that, unlike regression, where fitting the underlying data manifold reveals the distinct inductive biases of each model, classification operates on overlapping class distributions, where performance depends on both boundary placement and probability calibration. As a result, the mapping from data geometry to model performance becomes inherently less stable and highly dataset-specific, preventing the emergence of a consistent correspondence between dataset characteristics and algorithm behavior. Consequently, although models induce different decision functions and probability estimates, their inductive biases are not reliably reflected in observable performance patterns, even when evaluated using proper scoring rules such as the Brier score.
As a result, reliable a priori algorithm selection in classification remains fundamentally challenging.
6.3 Driving Factors of Dataset Difficulty
By extracting feature importances from our meta-models, we identified the specific data complexity measures that drive problem difficulty. For regression, our analysis highlights target correlation ($C1_{r}$), smoothness (S3), and non-linearity (L1) as the most critical determinants of model success. This perfectly corroborates the foundational regression complexity framework recently proposed by maciel_measuring_2016, who introduced these metrics to the regression domain. Specifically, lorena2018data adapted maximum feature correlation, nearest-neighbor error (smoothness), and linear error from classification to regression, and our findings confirm that these specific geometric properties, rather than basic dimensionality or sparsity metrics (N,d, T2), dictate the success of interpretable regression models. For classification, our meta-model relied heavily on the Error Rate of the Nearest Neighbor Classifier (N3) and the Maximum Fisher’s Discriminant Ratio (F1). These metrics are foundational classification complexity measures originally proposed by ho2002complexity and are widely recognized in modern meta-learning surveys as powerful descriptors of class overlap and decision boundary complexity [rivolli2018characterizing]. The prominence of F1 and N3 in our results confirms that the main source of difficulty in interpretable classification lies in the geometric interleaving of classes, rather than structural data properties [rivolli2018characterizing]. We must note a discrepancy in predictive strength, however: while these meta-features act as robust predictors of model success in regression, their influence is significantly less pronounced in classification scenarios.
6.4 Interpretability and Computational Tax
In interpretable machine learning, predictive accuracy must be weighed against structural complexity and computational efficiency. Measuring interpretability objectively is notoriously ambiguous, as noted by kruschel2025challenging, whose criteria yielded counter-intuitive results (e.g., ranking Decision Trees low on interpretability).
By comparing models within the same structural families, we observed clear trade-offs. Within linear models, extending the feature space (PR+LASSO) captures non-linear relationships and boosts accuracy but inherently sacrifices transparency by exploding the number terms compared to standard linear models. Similarly, for tree-based models, GOSDT produces significantly sparser and more interpretable trees than traditional Decision Trees, but this structural parsimony demands up to two orders of magnitude more computational time. More expressive methods (EBM, IGANN, PR+LASSO, SR) universally incur heavy computational penalties, highlighting that the search for optimal, inherently interpretable representations is fundamentally bounded by execution time constraints.
6.5 Practical Guidelines for Practitioners
Based on our analysis, we offer the following guidelines for selecting interpretable machine learning models. The decision process is heavily dependent on the specific type of task, the available computational budget, and the desired form of interpretability.
For Regression Tasks
The regression landscape is highly structured, allowing practitioners to make informed, a priori algorithm selections based on dataset characteristics.
  • Default to EBM or SR for Accuracy: EBM and Symbolic Regression consistently yield the highest predictive performance.
  • Account for Out-of-Sample Robustness: Models like Decision Trees and PR+LASSO are more sensitive and prone to performance degradation on unseen data, whereas strictly regularized models like LASSO provide superior out-of-sample stability.
  • Beware of Scalability Limits: While SR is highly competitive, its computationally expensive search process makes it impractical for very large datasets or strictly time-constrained environments.
  • Leverage Dataset Meta-Features: Practitioners should compute dataset meta-features to guide model selection (e.g., use LASSO over k-NN when dealing with highly sparse data, or rely on SR, GAMs, or Decision Trees rather than linear models when the data exhibits high non-linearity).
  • Weigh Expressivity Against Time: Models that capture complex, non-linear relationships (like PR+LASSO or IGANN) incur a massive computational penalty compared to standard linear models or decision trees.
For Classification Tasks
Unlike regression, classification performance is highly idiosyncratic and heavily dataset-dependent, making a priori algorithm selection difficult.
  • Adopt an Empirical Approach: Apart from a slight overall advantage for EBM, performance distributions overlap significantly. Because meta-features cannot reliably predict the best-performing model, practitioners must allocate time to train and evaluate multiple models.
  • Factor in the Interpretability Tax: For practitioners who strictly require a specific model structure (e.g., decision trees), there is a stark trade-off between structural simplicity and training time. For instance, while GOSDT produces significantly sparser and more interpretable trees than traditional Decision Trees, it requires up to two orders of magnitude more computational time.